# ACT-Tensor: Tensor Completion Framework for Financial Dataset Imputation

### v1

@@ 5 Experiment
Table 1: Setting of Parameters
@@
Table 2: Imputation Experiment Results
@@
Out-of-Sample MAR Out-of-Sample Block Out-of-Sample Logit
[…]
@@
5.1 Experiment Settings
Dataset. Our empirical analysis draws on the CRSP/Compustat monthly panel for 22,630 U.S. common stocks from January 2016 through December 2020. We apply standard filters to drop delisted stocks, extreme outliers, and non-exchange listings. For each firm’s monthly observation, we retain 45 widely used characteristics spanning fundamentals, momentum, sentiment, and trading frictions. To ensure all characteristics are on a comparable, unit-free scale, we cross-sectionally rank each characteristic every month, recenter the ranks to zero, and linearly rescale them to the interval [0.5,0.5][-0.5,0.5][ - 0.5 , 0.5 ]. Roughly 83% of the firm-characteristic entries are missing, highlighting the critical importance of a robust imputation strategy.
Benchmarks. We benchmark ACT-Tensor against both ad-hoc and state-of-the-art methods: (i) Cross-Sectional Median fills each missing data with the cross-sectional median of that characteristic; (ii) Global BF-XS [6] applies bidirectional temporal interpolation followed by a multi-factor ridge regression; and (iii) Local B-XS [6] uses only backward filling and a multi-factor ridge regression with a rolling window for factor estimation.
Missing Regimes. Since naturally missing entries lack ground truth values, we assess imputation accuracy by artificially introducing calibrated missingness: 10% of the fully observed cells are masked to create a held-out testset, which we then evaluate under three masking regimes: (i) Missing-at-random (MAR), where data is masked completely at random; (ii) Block Missingness (Block), which randomly mask data in contiguous one-year blocks where roughly 40% of these blocks are placed at the start of the sample, following [6]; and (iii) Logistic Missingness (Logit), where a two-stage logistic model first predicts whether a data begins with an initial gap and then assigns month-by-month missing probabilities conditional on past observations, producing heterogeneous, firm-specific patterns [6].
@@ 5.1 Experiment Settings
5.2 Overall Imputation Performance
@@
Large Gains under Structured Missingness. ACT-Tensor’s advantage is most significant when missingness follows a clear structure. In the Block regime—where entire one-year windows disappear—our model boosts $R^{2}_{\text{imp}}$ by 17% over the strongest baseline, and in the firm-dependent Logit pattern it nearly doubles $R^{2}_{\text{imp}}$ from 46% to 88%. These gains stem from ACT–Tensor’s ability to preserve the panel’s tensor structure, allowing it to recover entire blocks of data by exploiting cross-firm co-movements that matrix methods discard when they flatten the panel.
@@
Table 3: Asset Pricing Experiment Results
@@
5.3 Sparse-Cluster Stress Test
Imputation Stability under Extreme Sparsity. ACT-Tensor demonstrates exceptional stability under extreme sparsity conditions. When zooming in from the overall panel to its most challenging portion with extreme sparsity, we observe that over 80% of companies in the dataset have fewer than 3% of their data entries. In this demanding setting, as shown in panel B of Table 2: under the MAR regime, ACT-Tensor improves $R^{2}_{\text{imp}}$ by an exceptional 74.5% over the best benchmark. It achieves similarly large improvements in the Block and Logit scenarios, cutting the RMSE by approximately 20% and increasing R² by at least 45%, confirming its stability where other methods fail. The designed architecture of ACT-Tensor is uniquely suited for extreme sparsity, as it preserves the panel’s multi-way structure, allows sparse clusters to borrow statistical strength from dense ones, and applies temporal smoothing to ensure robust outputs.
5.4 Ablation Study
@@
Each Module Excels in Distinct Regimes. We isolate the cluster-based completion and temporal smoothing modules to assess their independent effectiveness. In our experiments, we compare each module against the vanilla CP completion with the same imputation settings. As shown in Panel C of Table 2, both modules improve upon the standard CP completion but excel under different missingness regimes. Temporal smoothing performs best in the MAR regime, where it leverages time-series continuity to handle sporadic noise, while cluster-based completion delivers the largest improvements in the Block and Logit settings by capturing cross-sectional heterogeneity.
Modules’ Effects Are Cumulative. We also tested the combined effects of the two modules under different missingness regimes, and the results confirm that their synergy drives the model’s superior performance. Temporal smoothing module excels when missingness is random, leveraging time-series continuity to handle sporadic noise. However, in cases of block-like or firm-specific missingness, smoothing alone can obscure important cross-sectional differences. In these scenarios, the cluster-based completion module first groups firms with similar data density, imputing within each cluster and preserving latent patterns. ACT-Tensor applies these steps in the following order: it first imputes within density-matched clusters, and then applies temporal smoothing to remove short-lived noise. This sequence ensures both stable cross-sectional structure and robust time-series trends, delivering the highest accuracy across all missingness regimes.
@@
5.5 Regularization-Free Stability Test
@@
Figure 1: RMSE Sensitivity under (a) Block and (b) Logistic Missing Regimes Against Regularization Coefficient $\lambda$.
We initially included an $\ell_{2}$ penalty in our model as a standard precaution to guard against potential overfitting, particularly when imputing the sparse clusters shown in Equation 11. To determine if this regularization was truly necessary, we conducted an experiment to test its impact on performance. We systematically varied the regularization weight, $\lambda$ across a wide range from very small $10^{-5}$ to moderate (0.50.50.5) and recorded the out-of-sample RMSE under the Block and Logit missingness regimes.
@@
5.6 Asset Pricing Performance Analysis
@@
6 Conclusion
In this paper, we propose ACT-Tensor, a robust and flexible tensor completion framework that tackles the pervasive, heterogeneous missing-data problem in asset pricing models. Our approach successfully overcomes the key limitations of conventional methods by preserving the dataset’s multidimensional structure while remaining robust to the extreme and heterogeneous patterns of missingness common in financial data. The strength of ACT-Tensor lies in its design of two innovative modules: cluster-based completion, which captures the heterogeneous cross-sectional patterns, and temporal smoothing, which filters out short-lived noise while preserving slow-moving fundamentals. Our extensive experiments demonstrated that this architecture delivers statistically superior imputations, with particularly strong performance in the most challenging sparse-cluster stress test. Crucially, we showed that this statistical accuracy translates directly into financial utility. Data imputed by ACT-Tensor led to the construction of accurate asset-pricing models and investment strategies that deliver higher risk-adjusted returns. This confirms that our framework not only fills in missing values with high precision but also preserves the essential financial signals required for effective quantitative analysis and decision-making.

### v3

@@ 5. Experiment
Table 1. Setting of Parameters
@@
Table 2. Imputation Experiment Results
@@
CP Completion 0.1511 0.0986 1.2894 0.6961 0.1695 0.1128 1.5100 0.6724 0.1980 0.1338 1.4922 0.7646
CP Completion w/ Clustering 0.1431 0.0922 1.2601 0.7292 0.1632 0.1070 1.4553 0.6958 0.1938 0.1293 1.4448 0.8016
CP Completion w/ CMA 0.1351 0.0782 1.0689 0.7585 0.1655 0.1085 1.4515 0.6511 0.1817 0.1160 1.2749 0.5637
ACT-Tensor w/ CMA 0.1321 0.0753 1.0522 0.7748 0.1599 0.1032 1.3941 0.7047 0.1803 0.1134 1.2489 0.8796
@@
5.1. Experiment Settings
Dataset. Our empirical analysis draws on the CRSP/Compustat monthly panel for 22,630 U.S. common stocks from January 2016 through December 2020. We apply standard filters to drop delisted stocks, extreme outliers, and non-exchange listings. For each firm’s monthly observation, we retain 45 widely used characteristics spanning fundamentals, momentum, sentiment, and trading frictions. To ensure all characteristics are on a comparable, unit‑free scale, we cross‑sectionally rank each characteristic every month, recenter the ranks to zero, and linearly rescale them to the interval [0.5,0.5][-0.5,0.5]. Roughly 83% of the firm-characteristic entries are missing, highlighting the critical importance of a robust imputation strategy.
Benchmarks. We benchmark ACT-Tensor against both ad-hoc and state-of-the-art methods: (i) Cross-Sectional Median fills each missing data with the cross‐sectional median of that characteristic; (ii) Global BF-XS (Bryzgalova et al., 2025) applies bidirectional temporal interpolation followed by a multi-factor ridge regression; and (iii) Local B-XS (Bryzgalova et al., 2025) uses only backward filling and a multi‐factor ridge regression with a rolling window for factor estimation.
Missing Regimes. Since naturally missing entries lack ground truth values, we assess imputation accuracy by artificially introducing calibrated missingness: 10% of the fully observed cells are masked to create a held-out testset, which we then evaluate under three masking regimes: (i) Missing-at-random (MAR), where data is masked completely at random; (ii) Block Missingness (Block), which randomly mask data in contiguous one-year blocks where roughly 40% of these blocks are placed at the start of the sample, following (Bryzgalova et al., 2025); and (iii) Logistic Missingness (Logit), where a two‑stage logistic model first predicts whether a data begins with an initial gap and then assigns month‑by‑month missing probabilities conditional on past observations, producing heterogeneous, firm‑specific patterns (Bryzgalova et al., 2025).
@@ 5.1. Experiment Settings
5.2. Overall Imputation Performance
@@
Large Gains under Structured Missingness. ACT-Tensor’s advantage is most significant when missingness follows a clear structure. In the Block regime—where entire one‑year windows disappear—our model boosts $R^{2}_{\text{imp}}$ by 17% over the strongest baseline, and in the firm‑dependent Logit pattern it nearly doubles $R^{2}_{\text{imp}}$ from 46% to 88%. These gains stem from ACT–Tensor’s ability to preserve the panel’s tensor structure, allowing it to recover entire blocks of data by exploiting cross‑firm co‑movements that matrix methods discard when they flatten the panel.
@@
Table 3. Asset Pricing Experiment Results
@@
5.3. Sparse-Cluster Stress Test
Imputation Stability under Extreme Sparsity. ACT-Tensor demonstrates exceptional stability under extreme sparsity conditions. When zooming in from the overall panel to its most challenging portion with extreme sparsity, we observe that over 80% of companies in the dataset have fewer than 10% of their data entries. In this demanding setting, as shown in panel B of Table 2: under the MAR regime, ACT-Tensor improves $R^{2}_{\text{imp}}$ by an exceptional 74.5% over the best benchmark. It achieves similarly large improvements in the Block and Logit scenarios, cutting the RMSE by approximately 20% and increasing R² by at least 45%, confirming its stability where other methods fail. The designed architecture of ACT-Tensor is uniquely suited for extreme sparsity, as it preserves the panel’s multi-way structure, allows sparse clusters to borrow statistical strength from dense ones, and applies temporal smoothing to ensure robust outputs.
5.4. Ablation Study
@@
Each Module Excels in Distinct Regimes. We isolate the cluster-based completion and temporal smoothing modules to assess their independent effectiveness. In our experiments, we compare each module against the vanilla CP completion with the same imputation settings. As shown in Panel C of Table 2, both modules improve upon the standard CP completion but excel under different missingness regimes. Temporal smoothing performs best in the MAR regime, where it leverages time‑series continuity to handle sporadic noise, while cluster-based completion delivers the largest improvements in the Block and Logit settings by capturing cross‑sectional heterogeneity.
Modules’ Effects Are Cumulative. We also tested the combined effects of the two modules under different missingness regimes, and the results confirm that their synergy drives the model’s superior performance. Temporal smoothing module excels when missingness is random, leveraging time‑series continuity to handle sporadic noise. However, in cases of block‑like or firm‑specific missingness, smoothing alone can obscure important cross‑sectional differences. In these scenarios, the cluster-based completion module first groups firms with similar data density, imputing within each cluster and preserving latent patterns. ACT‑Tensor applies these steps in the following order: it first imputes within density‑matched clusters, and then applies temporal smoothing to remove short‑lived noise. This sequence ensures both stable cross‑sectional structure and robust time‑series trends, delivering the highest accuracy across all missingness regimes.
@@
5.5. Regularization‑Free Stability Test
@@
Figure 1. RMSE Sensitivity under (a) Block and (b) Logistic Missing Regimes Against the Regularization Coefficient $\lambda$.
We initially included an $\ell_{2}$ penalty in our model as a standard precaution to guard against potential overfitting, particularly when imputing the sparse clusters shown in Equation 11. To determine if this regularization was truly necessary, we conducted an experiment to test its impact on performance. We systematically varied the regularization weight, $\lambda$ across a wide range from very small $10^{-5}$ to moderate ($0.5$) and recorded the out-of-sample RMSE under the Block and Logit missingness regimes.
@@
5.6. Asset Pricing Performance Analysis
@@
6. Conclusion
In this paper, we propose ACT-Tensor, a robust and flexible tensor completion framework that tackles the pervasive, heterogeneous missing‑data problem in asset pricing models. Our approach successfully overcomes the key limitations of conventional methods by preserving the dataset’s multidimensional structure while remaining robust to the extreme and heterogeneous patterns of missingness common in financial data. The strength of ACT-Tensor lies in its design of two innovative modules: cluster-based completion, which captures the heterogeneous cross-sectional patterns, and temporal smoothing, which filters out short‑lived noise while preserving slow‑moving fundamentals. Our extensive experiments demonstrated that this architecture delivers statistically superior imputations, with particularly strong performance in the most challenging sparse-cluster stress test. Crucially, we showed that this statistical accuracy translates directly into financial utility. Data imputed by ACT-Tensor led to the construction of accurate asset-pricing models and investment strategies that deliver higher risk-adjusted returns. This confirms that our framework not only fills in missing values with high precision but also preserves the essential financial signals required for effective quantitative analysis and decision-making.
Acknowledgements.
Elynn Chen’s research is supported in part by the NSF Award 2412577.